OpenStudy (anonymous):

given that the x intercept of a line is (1,0) and the y intercept is (0,-3) what is the eqaution of the line in slope intercept form?

OpenStudy (anonymous):

M=(0-(-3))/(1-0)=3 y-0=3(x-1) y=3x-3

OpenStudy (anonymous):

Is that the form you were looking for?

OpenStudy (anonymous):

yes it was thank you very much! :)

OpenStudy (anonymous):

My pleasure :).